A bullet of mass 20g is horizontally fired with a velocity 150 m s-1 from pistol of mass 2 kg. What is the recoil velocity of the pistol ?


1
Expert's answer
2021-02-22T10:20:05-0500

The momentum conservation law states, that the momentum of the bullet is equal and opposite (in direction) to the momentum of the pistol:


"p_b = p_p\Rightarrow m_bv_b = m_pv_p"

where "m_b = 20g = 0.02kg, m_p = 2kg" are masses, and "v_b = 150m\/s, v_p" and velocities of the bullet and the pistol respectively. Thus, expressing "v_p", obtain:


"v_p = \dfrac{m_bv_b}{m_p}\\
v_p = \dfrac{0.02kg\cdot 150m/s}{2kg} = 1.5 m/s"

Answer. 1.5 m/s.